OpenStudy (anonymous):

find the value of tan (sin ^-1(1/2))

OpenStudy (misty1212):

hi

OpenStudy (misty1212):

do you know a number whose sine is one half?

OpenStudy (anonymous):

hey

OpenStudy (misty1212):

or an "angle" if you want to think that way an angle whose sine is one half?

OpenStudy (anonymous):

those are the choices @misty1212

OpenStudy (misty1212):

no they are not

OpenStudy (misty1212):

your answer should have a \(\pi\) in it if you are working in radians

OpenStudy (vishweshshrimali5):

First, see: \[\large{\sin^{-1} x = y}\] means that: \[\large{x = \sin y}\]

OpenStudy (misty1212):

\[\sin(\theta)=\frac{1}{2}\] find \(\theta\) it is none of those, they must be answers to a different question

OpenStudy (misty1212):

oh dear, did you change the question on me?

OpenStudy (vishweshshrimali5):

@misty1212 You have to find \(\tan(\theta)\) :)

OpenStudy (anonymous):

lol this is the full question

OpenStudy (misty1212):

\[\tan(\arcsin(\frac{1}{2}))=\tan(\frac{\pi}{6})\] since \[\sin(\frac{\pi}{6})=\frac{1}{2}\]

OpenStudy (misty1212):

all you need now is \[\tan(\frac{\pi}{6})\]

OpenStudy (anonymous):

so choice a is the answer?

OpenStudy (misty1212):

no

OpenStudy (anonymous):

pi is 3

OpenStudy (vishweshshrimali5):

pi is 180 in degrees